# EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation

### 10 Mar 2026 v2

@@ 6 Conclusion
We presented EVM-QuestBench, an execution-grounded benchmark for natural language transaction script generation on EVM-compatible chains, instantiated on BNB Smart Chain (chain ID 56). The benchmark contains 107 tasks with Atomic and Composite splits. Evaluation executes model generated TypeScript scripts in a forked environment with snapshot isolation. Validators score outcomes after execution. Composite tasks apply a step efficiency factor.
We evaluated 20 models under a unified runner with 5 independent evaluation rounds per model. Performance varies substantially across models. Split scores expose a persistent gap between single transaction precision and workflow completion. Several models achieve strong Composite scores with weaker Atomic scores. Several models fail on Composite tasks despite non zero Atomic performance.
EVM-QuestBench offers a standardized protocol for studying execution-grounded behavior in on-chain automation. Future work can expand task coverage, extend the benchmark to additional EVM chains, and strengthen analysis using per task traces. In fact, we have already leveraged the same architecture to develop and establish a benchmark on Solana, demonstrating the portability of our atomic/composite paradigm across heterogeneous blockchain ecosystems. We also plan to incorporate richer security checks for transaction intent and side effects. Additionally, we intend to evaluate LLMs’ success rates in generating atomic and composite task definitions, which would further assess their understanding of on-chain transactions and smart contract semantics.
@@ 7 Limitations
Because EVM-QuestBench is execution-grounded, scores inherit instability from RPC connectivity, fork performance, and provider availability. Snapshot isolation and composite retries reduce, but do not eliminate, this effect.
Reproducibility across forks.
By default, the fork block height is not pinned, and cross-run results can drift with mainnet state, especially for price-sensitive swaps and liquidity operations. Our framework supports pinning the fork at a specific block height via a configuration file; we recommend this setting for standardized comparative benchmarking. Recording the fork block number in evaluation metadata further aids reproducibility.
Scoring and interface effects.
Composite scoring applies a step efficiency decay factor $\min(1,K_{\mathrm{opt}}/K_{\mathrm{act}})$, where $K_{\mathrm{act}}$ counts execution rounds that may include retries; this can conflate planning inefficiency with execution instability. Validators also use tolerances for some tasks, which may admit occasional false positives or over penalize benign deviations. Finally, providers differ in default output budgets and formatting behavior, which can affect schema validity; prior work has noted that automatic metrics and LLM as a judge can introduce additional biases (zheng2023judging; liu2023geval).
@@
We conduct 5 independent evaluation rounds per model and report mean scores with confidence intervals (Table 1). While 5 rounds provide reasonable statistical power for tier-level separation, some adjacent-rank differences remain statistically indistinguishable (overlapping CIs). Additional rounds would narrow confidence intervals and enable finer-grained ranking distinctions.
Appendix A Reproducibility
@@ A.3 Key Experimental Settings
Runtime TypeScript runner using ethers.js v6
@@
Appendix B Prompt and Interface Specification
@@ B.1 Runner Interface
B.2 Atomic and Composite Prompt Roles (Schematic)
@@
B.3 Schema Invalid Rules
@@
Appendix C Task Definition Schema
@@
C.1 Atomic Task Fields
@@
C.2 Composite Task Fields
@@
Appendix D Validators and Scoring
@@ D.1 Validator Families
Native and ERC20 transfer small relative tolerance receipt success, recipient, amount, balance delta
Approval exact match spender address, allowance target
@@
D.2 Composite Step Efficiency Decay
@@
D.3 Validator Tolerance Sensitivity
@@
Appendix E Additional Analysis Figures
@@
E.1 Total Score Ranking
@@
E.2 Composite Pattern Difficulty
@@
E.3 Step Overhead Distribution
@@
E.4 Atomic Subcategory Difficulty
@@
Appendix F Full Evaluation Tables
@@ F.1 Quadrant Summary
F.2 Model Coverage and Quadrant Classification
@@
F.3 Step Overhead and Decay Impact Summary
@@ Appendix G End-to-End Walkthrough
    Balance Change (30 pts): sender balance decreased by amount ++ gas; recipient balance increased by amount, both within 1% tolerance. ✓
@@
The expected transfer amount is computed dynamically from the fork state (not hardcoded), so the ground truth is always consistent with the execution environment. The tolerance on balance change (1%) absorbs gas cost estimation differences without admitting incorrect transfers.
Appendix H Case Studies
@@
H.1 Case 1: Complex Multi step DeFi Workflow
@@
H.2 Case 2: Batch Approval Workflow
@@
H.3 Case 3: Natural Language Misinterpretation
@@
H.4 Case 4: Planning-Level Failure
@@
H.5 Case 5: Partial Execution in Multi-Step Workflows

### 16 Mar 2026 v3

@@ 6 Conclusion
We presented EVM-QuestBench, an execution-grounded benchmark for natural language transaction script generation on EVM-compatible chains, instantiated on BNB Smart Chain (chain ID 56). The benchmark contains 107 tasks with Atomic and Composite splits. Model-generated TypeScript scripts are executed in a snapshot-isolated forked environment and scored by validators against post-state constraints; composite tasks additionally apply a step efficiency factor.
We evaluated 20 models under a unified runner with 5 independent rounds each. Results reveal a persistent capability gap between single-transaction precision and multi-step workflow completion. Thanks to its dynamic parameterization design, a full evaluation run requires as few as 334K tokens (as low as \$0.06), making large-scale, multi-round model comparison practically affordable.
EVM-QuestBench offers a standardized protocol for studying execution-grounded behavior in on-chain automation. We have already ported the same architecture to Solana, demonstrating the portability of the atomic/composite paradigm across heterogeneous blockchain ecosystems. Future work will expand task coverage, incorporate richer security checks for transaction intent and side effects, and evaluate LLMs’ ability to generate task definitions autonomously.
@@ 7 Limitations
Because EVM-QuestBench is execution-grounded, scores inherit potential instability from RPC connectivity, fork performance, and provider availability. Although this effect was negligible in our experiments, snapshot isolation and composite retries reduce but cannot fully eliminate such instability.
@@
We conduct 5 independent evaluation rounds per model and report mean scores with standard deviations and observed ranges (Table 1). While 5 rounds provide reasonable statistical power for tier-level separation, some adjacent-rank differences remain difficult to distinguish. Additional rounds would provide more stable variance estimates and enable finer-grained ranking distinctions.
Task coverage.
The current benchmark contains 107 tasks, which covers a representative range of on-chain operations but remains limited in scale. We encourage community contributions to expand task coverage across protocols, chains, and operation types, and plan to maintain EVM-QuestBench as a living benchmark with ongoing community participation.
Appendix A Background: EVM Concepts and What EVM-QuestBench Measures
This section provides a self-contained introduction to key concepts from the EVM ecosystem for readers unfamiliar with blockchain technology, and explains concretely what EVM-QuestBench is designed to evaluate.
A.1 Core EVM Terminology
Ethereum Virtual Machine (EVM).
The EVM is a sandboxed, stack-based virtual machine that executes smart contract bytecode on a distributed network of nodes. Every EVM-compatible chain—including Ethereum mainnet, BNB Smart Chain (BSC), Polygon, and Avalanche—runs the same instruction set, so programs written for one chain can generally be ported to others with minimal changes. EVM-QuestBench is instantiated on BSC (chain ID 56) but the design generalizes to any EVM-compatible chain.
Smart Contract.
A smart contract is a piece of code deployed at a fixed address on the blockchain. Once deployed, it can be called by anyone by sending a transaction that encodes the function name and its arguments as raw bytes (calldata). The contract’s logic executes deterministically on every node, and the resulting state changes (e.g., token balances, ownership records) are permanently recorded on-chain. Errors in calldata construction or parameter encoding commonly cause transactions to revert.
Externally Owned Account (EOA).
An EOA is a user-controlled wallet identified by a public/private key pair. All transactions must originate from an EOA, which signs the transaction with its private key to authorize it. EVM-QuestBench creates a fresh test EOA for each evaluation run; the model generates unsigned transaction payloads that the runner signs and submits.
Transaction and Gas.
A transaction is an authenticated message that transfers value or invokes a smart contract. Each transaction consumes gas, a unit of computational cost paid by the sender in native tokens (BNB on BSC). Gas limits prevent infinite loops and prioritize efficient code. Incorrectly estimated gas limits can cause transactions to revert or fail silently.
BNB Smart Chain (BSC).
BSC is an EVM-compatible blockchain operated by Binance with fast block times (${\sim}3$s) and low fees. It supports the same tooling as Ethereum. EVM-QuestBench forks BSC mainnet via Anvil to provide a realistic yet isolated on-chain environment.
Token Standards.
  • ERC-20: Fungible token standard. Each contract maintains a balance mapping; transfers are authorized via approve/transferFrom or direct transfer. Amounts must be expressed in the token’s native unit (e.g., $10^{18}$ for 18-decimal tokens).
  • ERC-721: Non-fungible token (NFT) standard. Each token ID is unique and owned by exactly one address.
  • ERC-1155: Multi-token standard that supports both fungible and non-fungible assets in a single contract.
DeFi, DEX, and AMM.
Decentralized Finance (DeFi) refers to financial services built as smart contracts. A Decentralized Exchange (DEX) allows token swaps without a centralized intermediary. Most DEXes use an Automated Market Maker (AMM) model, where liquidity providers deposit token pairs into pools, and prices are determined by the ratio of reserves. PancakeSwap—the primary DEX used in EVM-QuestBench—is an AMM DEX on BSC. Swap transactions must encode the trade path, deadline, and minimum output amount; incorrect parameters cause reverts or financial loss.
Liquidity Provision and Staking.
Liquidity providers deposit two tokens into an AMM pool and receive LP tokens representing their share. LP tokens can be deposited into staking contracts to earn rewards. These workflows require multiple sequential transactions (approve $\to$ add liquidity $\to$ stake), making them natural candidates for composite task evaluation.
A.2 What EVM-QuestBench Measures
EVM-QuestBench evaluates whether a language model can translate a natural language instruction (e.g., “Swap 0.1 BNB for USDT and stake the LP tokens”) into a correct, executable client-side TypeScript module. The module must:
  1. 1.
    Correctly identify the target contract and function. The model must select the right protocol (e.g., PancakeSwap Router vs. staking pool), the right function signature, and encode calldata that satisfies the ABI.
  2. 2.
    Handle chain-specific units. Token amounts must be converted from human-readable form to on-chain representation (e.g., $0.1\times 10^{18}$ wei for 18-decimal tokens). Swaps require slippage-tolerant minimum output values.
  3. 3.
    Satisfy protocol prerequisites. Many operations require a prior approval transaction (ERC-20 approve) before the main action can execute. The model must identify and include these dependencies.
  4. 4.
    Propagate parameters across steps. In multi-step workflows, outputs from earlier steps (e.g., LP token amounts received from liquidity addition) feed into subsequent steps (e.g., staking). The model must track and propagate these values correctly.
The benchmark does not evaluate contract deployment or Solidity code generation. It specifically targets the client-side scripting layer: producing unsigned transaction payloads that a standardized runner can sign and broadcast. Validators then check the resulting on-chain state—not the code itself—ensuring that evaluation measures functional correctness under execution.
Appendix B Figures
This section contains benchmark composition and result figures referenced from the main text.
B.1 Benchmark Composition
B.2 Results Figures
Appendix C Reproducibility
@@ C.3 Key Experimental Settings
Runtime TypeScript runner using ethers.js v6 (primary; viem v2 also supported)
@@
Appendix D Prompt and Interface Specification
@@ D.1 Runner Interface
D.2 Atomic and Composite Prompt Roles (Schematic)
@@
D.3 Schema Invalid Rules
@@
Appendix E Task Definition Schema
@@
E.1 Atomic Task Fields
@@
E.2 Composite Task Fields
@@
Appendix F NL Template Difficulty Scoring
To control for potential template-wording bias, we implemented a difficulty scoring system for all 373 natural language templates. Three SOTA models (Claude Opus 4.6, GPT-5.4, and Gemini 3.1 Pro) independently rated each template on a 1–5 clarity scale (1 = maximal precision, 5 = maximal ambiguity), and ratings were averaged. Templates are classified into three tiers: precise ($\leq 2.0$), moderate (2.02.03.53.5), and vague (>3.5>3.5). Scores are stored in nl_template_scores.json.
The distribution is: 336 precise (90.1%), 37 moderate (9.9%), and 0 vague. This confirms that the template pool is consistently unambiguous across all task types.
At evaluation time, the --nl-difficulty flag supports four selection modes: random (default, no filtering), precise, moderate, and vague. All experiments in this paper use random mode. The difficulty API is available for controlled ablation studies on instruction clarity.
Appendix G Validators and Scoring
@@ G.1 Validator Families
Native and ERC20 transfer 0.1% relative receipt success, recipient, amount, balance delta
Approval exact match (====) spender address, allowance target
@@
G.2 Composite Step Efficiency Decay
@@
G.3 Validator Tolerance Sensitivity
@@
Appendix H Additional Analysis Figures
@@
H.1 Total Score Ranking
@@
H.2 Where Workflows Fail
Figure 9 reports pass rates by composite workflow pattern. The hardest patterns are multi-stage DeFi workflows combining liquidity and staking, as well as batch approval patterns. These concentrate three failure surfaces: (i) prerequisite correctness—the model must issue all required approvals before a protocol action; (ii) cross-step parameter consistency—output values from earlier steps (e.g., LP tokens received) must be correctly propagated to later steps; and (iii) execution robustness under multiple sequential transactions, where any single revert terminates the workflow and yields a zero base score.
H.3 Composite Pattern Difficulty
@@
H.4 Step Overhead Distribution
@@
H.5 Atomic Subcategory Difficulty
@@
Appendix I Full Evaluation Tables
@@ I.1 Quadrant Summary
I.2 Model Coverage and Quadrant Classification
@@
I.3 Step Overhead and Decay Impact Summary
@@
    Balance Change (30 pts): sender balance decreased by amount ++ gas; recipient balance increased by amount, both within 0.1% tolerance. ✓
@@
The expected transfer amount is computed dynamically from the fork state (not hardcoded), so the ground truth is always consistent with the execution environment. The tolerance on balance change (0.1%) absorbs floating-point precision differences without admitting materially incorrect transfers.
Appendix K Case Studies
@@
K.1 Case 1: Complex Multi step DeFi Workflow
@@
K.2 Case 2: Batch Approval Workflow
@@
K.3 Case 3: Natural Language Misinterpretation
@@
K.4 Case 4: Planning-Level Failure
@@
K.5 Case 5: Partial Execution in Multi-Step Workflows
@@
Appendix L Token Usage and Cost Analysis
Figure 3 (Section 5) plots total token consumption against benchmark score for a single 107-task run. This appendix provides the detailed breakdown.
Token-efficient models.
Token budgets vary by over $4\times$ across the top-11 models. Claude-Sonnet-4.5 achieves the highest score (8,236) while consuming only 431K tokens per run, making it both the most accurate and one of the most token-efficient models. Other efficient models include DeepSeek-V3.2 (334K tokens, \$0.07), GPT-5.1 (358K, \$1.15), and Gemini-2.5-Flash (418K, \$0.29), all scoring above 6,700 with fewer than 420K tokens per run.
Thinking-enabled models.
Thinking-enabled models allocate substantially more output tokens to chain-of-thought reasoning: Kimi-K2-Thinking (1,636K, \$2.74), Gemini-3-Pro (1,496K, \$14.16), and GPT-5 (986K, \$7.48) each exceed 900K tokens, with output tokens dominating their budgets (82–93% of total).
Cost range and caching.
Cost spans three orders of magnitude. GPT-OSS-120B achieves rank 7 at just \$0.06 per run, making it $230\times$ cheaper than Gemini-3-Pro (rank 2, \$14.16) despite only a 12% score gap. Claude-Sonnet-4.5 incurs \$1.91 per run due to its premium output pricing (\$15/M), yet its low token count keeps it far below the thinking models in absolute cost. Prompt caching reduces input costs by up to 30% for models with high cache-read ratios (Claude, Gemini), though output-heavy thinking models benefit minimally (2–4% savings) since caching applies only to input tokens.